# Provably Unlearnable Examples

### v1

@@ Abstract.
The exploitation of publicly accessible data has led to escalating concerns regarding data privacy and intellectual property (IP) breaches in the age of artificial intelligence. As a strategy to safeguard both data privacy and IP-related domain knowledge, efforts have been undertaken to render shared data unlearnable for unauthorized models in the wild. Existing methods apply empirically optimized perturbations to the data in the hope of disrupting the correlation between the inputs and the corresponding labels such that the data samples are converted into Unlearnable Examples (UEs). Nevertheless, the absence of mechanisms that can verify how robust the UEs are against unknown unauthorized models and train-time techniques engenders several problems. First, the empirically optimized perturbations may suffer from the problem of cross-model generalization, which echoes the fact that the unauthorized models are usually unknown to the defender. Second, UEs can be mitigated by train-time techniques such as data augmentation and adversarial training. Furthermore, we find that a simple recovery attack can restore the clean-task performance of the classifiers trained on UEs by slightly perturbing the learned weights. To mitigate the aforementioned problems, in this paper, we propose a mechanism for certifying the so-called $(q,\eta)$-Learnability of an unlearnable dataset via parametric smoothing. A lower certified $(q,\eta)$-Learnability indicates a more robust protection over the dataset. Finally, we try to 1) improve the tightness of certified $(q,\eta)$-Learnability and 2) design Provably Unlearnable Examples (PUEs) which have reduced $(q,\eta)$-Learnability. According to experimental results, PUEs demonstrate both decreased certified $(q,\eta)$-Learnability and enhanced empirical robustness compared to existing UEs. Compared to the competitors on classifiers with uncertainty in parameters, PUEs reduce at most $18.9\%$ of certified $(q,\eta)$-Learnability on ImageNet and $54.4\%$ of the empirical test accuracy score on CIFAR-100.
@@
Figure 1. Recovery attacks using a small portion ($1\%$-$20\%$) of the CIFAR10 training set. Points on the curves trace the clean testing accuracy of classifiers whose weights are perturbed away from the poisoned classifier. For each fine-tuned classifier, the $\ell_{2}$ norm of the weight perturbation is capped by $\eta$. This special adversary reveals that current UEs are not robust against uncertainty in classifier parameters, and their reliability cannot be guaranteed. Therefore, a mechanism for certifying UE performance is pivotal.
1. Introduction
@@
Data privacy and intellectual property (IP) breaches have become major concerns in recent legislations and regulations serving for building responsible Artificial Intelligence (AI), such as GDPR (European Parliament and Council of the European Union, 5 04), CCPA (Legislature, 2018), the European Union AI Act (Commission, 2021), the recent US Executive Order on the Safe, Secure, and Trustworthy Development and Use of AI (House, 2023). These regulations have emphasized that consumers have the right to limit the use and disclosure of data collected from them. Despite the laws and regulations, technologies for enhancing data privacy and protecting IP face increasingly severe challenges due to rapidly evolving machine learning algorithms. Particularly, data published by individual users on content-sharing social media platforms is exposed to the threat of domain exploitation by unauthorized machine learning models. For instance, a pretrained Stable Diffusion can be fine-tuned on a small set of paintings to generate new images mimicking the style of the paintings (Heikkilä, 2022; Gal et al., 2022). As another example, classifiers trained on publicly accessible data can be used to launch membership inference attacks exposing confidential personal information (Shokri et al., 2017). As data becomes the new oil, these threats highlight the imminent need to implement effective measures for Data Availability Control (DAC).
@@
As a response, a series of attempts have been made to make the published data unlearnable. One pragmatic and effective way to protect image datasets against unauthorized classifiers is to perturb the pixels before publishing the images. Existing defenses apply anti-learning perturbations that minimize the training loss towards either correct labels or target labels to the input data (Fowl et al., 2021; Huang et al., 2021). In lieu of learning functions mapping the input data to the corresponding labels, the defenses encourage unauthorized models to learn a strong correlation (i.e., a shortcut) between the perturbations and the labels. As a consequence, despite that the trained models can achieve low error rates on the data with the anti-learning perturbations, they generalize poorly to datasets from the clean data distribution. This line of work is also referred to as Perturbative Availability Poison (PAP) in the literature (Feng et al., 2019; Tao et al., 2021; Fowl et al., 2021; Liu et al., 2023).
@@
Motivation.  Existing PAP methods empirically search for the anti-learning perturbations based on a finite amount of training samples and models (Huang et al., 2021; Fu et al., 2022; Tao et al., 2021; Fowl et al., 2021; Yuan and Wu, 2021; Chen et al., 2023; He et al., 2023). Henceforth, the generated perturbations may face the cross-model generalization problem brought by unknown illegal classifiers in the wild. Moreover, it is discovered that Unlearnable Examples (UEs) crafted by these PAP methods can be made invalid by prevailing train-time techniques such as data augmentations, adversarial training, and data compression algorithms (Huang et al., 2021; Fu et al., 2022; Liu et al., 2023). Even making fair comparisons of UEs is difficult due to the randomness, such as initialization and mini-batch selection, involved in the training process. In addition to these drawbacks, let us consider a type of recovery attack in which an attacker can fine-tune a poisoned classifier on a small set of clean data to restore its clean-task performance without significantly changing the classifier parameters. This kind of adversary may exist in federated learning systems based on weight aggregation with Byzantine robustness (Zhu et al., 2023). For example, an adversarial client training its unauthorized classifier on UEs can perform a recovery attack and upload the parameters to the server without alerting outlier detectors, rendering the global classifier to violate corresponding DAC protocols. We randomly collect $1\%$-$20\%$ of CIFAR10 training samples to fine-tune a ResNet-18 trained on the error-minimizing PAP (Huang et al., 2021). We use projected stochastic gradient descent (SGD) in the fine-tuning so the $\ell_{2}$ norm of the weights differences is restricted within a small range. Surprisingly, we find that the accuracy can be restored from around $0.1$ to near $0.8$ by seemingly minuscule modifications in the weights using $5\%$-$10\%$ of CIFAR10 samples (Figure 1). This result reflects that the robustness of UEs towards perturbations and uncertainties in classifier parameters is neither rigidly gauged nor carefully pondered.
@@
The aforementioned problems hinder the assessment of UEs and constitute additional risks to be exploited by deliberate adaptive adversaries. Therefore, it is crucial to develop a mechanism to ensure that a set of UEs can take effect without being ravaged by the problems of cross-model generalization and adaptive attackers. The mechanism can resort to finding classifiers that can achieve the best possible clean-task accuracy when trained on UEs and restraining this best-case accuracy. We are thus motivated to ask the question:
@@
[backgroundcolor=grey!10,rightline=true,leftline=true,topline=true,bottomline=true,roundcorner=1mm,everyline=true,nobreak=false] Is it possible to go beyond heuristics and provide an upper bound towards the clean-data performance of undivulged models trained on a set of UEs? The question is daunting to be thoroughly answered. However, we will show in this work that, it is possible to derive a clean-data accuracy upper bound guaranteed with a high probability, for unauthorized classifiers under some constraints. On the other hand, this threshold also serves as an upper bound of clean accuracy when facing the recovery attack. It can henceforth be a tool for better gauging the effectiveness and robustness of UEs.
Our method and its broader impact.  In this paper, we propose a mechanism for certifying the best clean-task accuracy retained by unauthorized classifiers trained on a set of UEs. The mechanism can be applied to any empirically made UEs to derive a certified $(q,\eta)$-Learnability. $(q,\eta)$-Learnability gauges the effectiveness and robustness of UEs towards arbitrary classifiers sampled from a certain parameter space. Note that the certification pertains to the parameter space rather than the input space, alleviating the need for knowledge of the training strategies or learning algorithms used by the attackers. We also proposed certifying a generalization $(q,\eta)$-Learnability, which removes the requirement for holding private test datasets. To make the certification more sound, the paper also proposes a simple method to expand the certifiable parameter space and reduce the gap between the certified $(q,\eta)$-Learnability and the True Learnability (see Definition 3.1). Moreover, we produce Provably Unlearnable Examples (PUEs) which achieve lower $(q,\eta)$-Learnability scores and be more robust than existing PAP methods.
The $(q,\eta)$-Learnability provides a rigid guarantee of data availability for machine learning models. This guarantee can operate in parallel with other techniques, such as deep watermarking (Wang et al., 2024) and differential privacy (DP) (Hu et al., 2023), to deter illegal data exploitation and support responsible AI regulations and standards (e.g., CCPA, GDPR, and Australia’s AI Safety Standard (Australia, 2023)). Instead of post-event ownership auditing through watermarks and protection of sensitive information with DP, $(q,\eta)$-Learnability proactively controls access to domain knowledge through data, safeguarding both data IP and privacy. The contributions of this paper are as follows:
@@
2. Background
@@ 2.1. Unlearnable Examples
(2) $\min_{\hat{f}}$ $\mathbb{E}_{(x,y)\sim{\mathbb{D}}_{s}}\min_{\delta}{\mathcal{L}}(\hat{f}(x+\delta),\,y)$
@@ 2.2. Smoothed Classifiers
(3) $g(x)=$ $\operatorname*{arg\,max}_{y\in{\mathcal{Y}}}{\Pr_{\epsilon\sim\pi(x)}[f_{\theta}(x+\epsilon)=y]},\hskip 10.00002ptor$
@@
$\pi(x)$ is the distribution of the noise (i.e., smoothing distribution) centered at $x$ (or $\theta$) and $y$ is a predicted label from the label space ${\mathcal{Y}}$. The output from the smoothed classifier can be statistically bounded to offer certifiable robustness with respect to changes in the smoothed variables. Cohen et al. derived a tight $\ell_{2}$ robustness bound by smoothing classifier inputs with Gaussian distributions and certifying based on Neyman-Pearson Lemma (Cohen et al., 2019). Subsequently, the certified robustness is extended to various smoothing noises and certification frameworks (Fischer et al., 2020; Li et al., 2021; Hao et al., 2022; sukenik2022intriguing; Cullen et al., 2022; Dvijotham et al., 2020; Zhang et al., 2020; Salman et al., 2020; Yang et al., 2020). Such smoothed classifier $g(x)$ can also be extended to randomized learning functions which in addition smooth the training dataset ${\mathbb{D}}_{s}$, resulting
(4) $g({\mathbb{D}}_{s},x)=\operatorname*{arg\,max}_{y\in{\mathcal{Y}}}\Pr_{\hat{{\mathbb{D}}}_{s}\sim\pi({\mathbb{D}}_{s})}[f_{\theta}(\hat{{\mathbb{D}}}_{s},\ x)=y].$
@@
Figure 2. An overview of the certification and PUE crafting framework. A dataset ${\mathbb{D}}_{s}$ is perturbed into ${\mathbb{D}}_{s}\oplus\delta$ before being released to the public. The $(q,\eta)$-Learnability of ${\mathbb{D}}_{s}\oplus\delta$ can be certified to ensure that any unauthorized classifier trained on ${\mathbb{D}}_{s}\oplus\delta$ has a provable upper bound on performance on any dataset ${\mathbb{D}}$ in the same domain with ${\mathbb{D}}_{s}\oplus\delta$, as long as the parameters of the unauthorized classifier are within a certified parameter set. Optimized PUEs lead to lower $(q,\eta)$-Learnability.
3. Problem Statement
Problem definition.  Inheriting from the previous definitions, consider a to-be-published dataset ${\mathbb{D}}_{s}$ (e.g., portraits to be posted on Facebook) with $N$ data points. Additionally, there is a private dataset ${\mathbb{D}}$ (e.g., patient records with photos in a hospital database) sharing the same distribution ${\mathcal{D}}$ with ${\mathbb{D}}_{s}$. The defender’s goal is to render ${\mathbb{D}}_{s}$ unlearnable against unknown learning algorithms and classifiers (e.g., deep neural networks) by perturbing it into ${\mathbb{D}}_{s}\oplus\delta$, where $\delta$ is the anti-learning perturbation and $\oplus$ is the perturbing operator. We specifically consider additive perturbations $\delta=\{\delta_{i}\}^{|{\mathbb{D}}_{s}|}_{i=1}$ applied to each sample of ${\mathbb{D}}_{s}$ in this paper. Thus, $\oplus$ represents the element-wise addition. “Unlearnable” here means that a classifier trained on ${\mathbb{D}}_{s}\oplus\delta$ would exhibit a high classification error on ${\mathbb{D}}$. Furthermore, the defender wants to verify how robust ${\mathbb{D}}_{s}\oplus\delta$ is by having an upper bound on the accuracy that can be achieved by any unknown classifier trained on ${\mathbb{D}}_{s}\oplus\delta$. More profoundly, the defender wants to determine the best possible accuracy that classifiers trained on ${\mathbb{D}}_{s}\oplus\delta$ can achieve on ${\mathbb{D}}$. The defender does not desire an elevated best possible accuracy since it implies ${\mathbb{D}}_{s}\oplus\delta$ is exploitable. Therefore, the research goals are as follows. {mdframed}[backgroundcolor=grey!10,rightline=true,leftline=true,topline=true,bottomline=true,roundcorner=1mm,everyline=true,nobreak=false]
  • A certification mechanism should be developed to verify the maximum accuracy achievable by any classifiers and learning algorithms on $\mathcal{D}$ through training on $\mathcal{D}_{s}\oplus\delta$.
  • A better noise $\delta$ should be designed such that $\mathcal{D}_{s}\oplus\delta$ can effectively decrease the verified best accuracy.

### v2

# Provably Unlearnable Data Examples

@@ Abstract
The exploitation of publicly accessible data has led to escalating concerns regarding data privacy and intellectual property (IP) breaches in the age of artificial intelligence. To safeguard both data privacy and IP-related domain knowledge, efforts have been undertaken to render shared data unlearnable for unauthorized models in the wild. Existing methods apply empirically optimized perturbations to the data in the hope of disrupting the correlation between the inputs and the corresponding labels such that the data samples are converted into Unlearnable Examples (UEs). Nevertheless, the absence of mechanisms to verify the robustness of UEs against uncertainty in unauthorized models and their training procedures engenders several under-explored challenges. First, it is hard to quantify the unlearnability of UEs against unauthorized adversaries from different runs of training, leaving the soundness of the defense in obscurity. Particularly, as a prevailing evaluation metric, empirical test accuracy faces generalization errors and may not plausibly represent the quality of UEs. This also leaves room for attackers, as there is no rigid guarantee of the maximal test accuracy achievable by attackers. Furthermore, we find that a simple recovery attack can restore the clean-task performance of the classifiers trained on UEs by slightly perturbing the learned weights. To mitigate the aforementioned problems, in this paper, we propose a mechanism for certifying the so-called $(q,\eta)$-Learnability of an unlearnable dataset via parametric smoothing. A lower certified $(q,\eta)$-Learnability indicates a more robust and effective protection over the dataset. Concretely, we 1) improve the tightness of certified $(q,\eta)$-Learnability and 2) design Provably Unlearnable Examples (PUEs) which have reduced $(q,\eta)$-Learnability. According to experimental results, PUEs demonstrate both decreased certified $(q,\eta)$-Learnability and enhanced empirical robustness compared to existing UEs. Compared to the competitors on classifiers with uncertainty in parameters, PUEs reduce at most $18.9\%$ of certified $(q,\eta)$-Learnability on ImageNet and $54.4\%$ of the empirical test accuracy score on CIFAR-100. Our source code is available at https://github.com/NeuralSec/certified-data-learnability.
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2025 24 - 28 February 2025, San Diego, CA, USA ISBN 979-8-9894372-8-3 https://dx.doi.org/10.14722/ndss.2025.24886 www.ndss-symposium.org
@@
Figure 1: Recovery attacks using a small portion ($1\%$-$20\%$) of the CIFAR10 training set. Points on the curves trace the clean testing accuracy of classifiers whose weights are perturbed away from the poisoned classifier. For each fine-tuned classifier, the $\ell_{2}$ norm of the weight perturbation is capped by $\eta$. This special adversary reveals that current UEs are not robust against uncertainty in classifier parameters, and their reliability cannot be guaranteed. Therefore, a mechanism for certifying UE performance is pivotal.
I Introduction
@@
Data privacy and intellectual property (IP) breaches have become major concerns in recent legislations and regulations serving for building responsible Artificial Intelligence (AI), such as GDPR [1], CCPA [2], the European Union AI Act [3], and the recent US Executive Order on the Safe, Secure, and Trustworthy Development and Use of AI [4]. These regulations have emphasized that consumers have the right to limit the use and disclosure of data collected from them. Despite the laws and regulations, technologies for enhancing data privacy and protecting IP face increasingly severe challenges due to rapidly evolving machine learning algorithms. Particularly, data published by individual users on content-sharing social media platforms is exposed to the threat of domain exploitation by unauthorized machine learning models. For instance, a pretrained Stable Diffusion can be fine-tuned on a small set of paintings to generate new images mimicking the style of the paintings [5, 6]. As another example, classifiers trained on publicly accessible data can be used to launch membership inference attacks exposing confidential personal information [7]. As data becomes the new oil, these threats highlight the imminent need to implement effective measures for Data Availability Control (DAC).
@@
As a response, a series of attempts have been made to make the published data unlearnable. One pragmatic and effective way to protect image datasets against unauthorized classifiers is to perturb the pixels before publishing the images. Existing defenses apply anti-learning perturbations that minimize the training loss with respect to either correct or target labels on the input data. [8, 9]. In lieu of learning functions mapping the input data to the corresponding labels, the defenses encourage unauthorized models to learn a strong correlation (i.e., a shortcut) between the perturbations and the labels. As a consequence, despite that the trained models can achieve low error rates on the data with the anti-learning perturbations, they generalize poorly to datasets from the clean data distribution. This line of work is also referred to as Perturbative Availability Poison (PAP) in the literature [10, 11, 8, 12].
@@
Motivation.  Existing PAP methods empirically search for the anti-learning perturbations based on a finite amount of training samples and models [9, 13, 11, 8, 14, 15, 16]. Henceforth, the generated perturbations may face the cross-model generalization problem brought by unknown illegal classifiers in the wild. Moreover, there is no rigid guarantee of the maximally attainable learning results for adversaries on the UEs. The widely adopted empirical test accuracy, as a metric, is prone to uncertainty in the data, model selection, and training process. This indicates that empirical test accuracy is not a sound metric for evaluating UEs. UEs resulting in lower test accuracy scores on certain models and datasets do not necessarily evince their superiority over others and may leave room for attackers. Let us consider a type of recovery attack in which an attacker can fine-tune a poisoned classifier on a small set of clean data to restore its clean-task performance without significantly changing the classifier parameters. This kind of adversary may exist in federated learning systems based on weight aggregation with Byzantine robustness [17]. For example, an adversarial client training its unauthorized classifier on UEs can perform a recovery attack and upload the parameters to the server without alerting outlier detectors, rendering the global classifier to violate corresponding DAC protocols. We randomly collect $1\%$-$20\%$ of CIFAR10 training samples to fine-tune a ResNet-18 trained on the error-minimizing PAP [9]. We use projected stochastic gradient descent (SGD) in the fine-tuning so the $\ell_{2}$ norm of the weight differences is restricted within a small range. Surprisingly, we find that the accuracy can be restored from around $0.1$ to near $0.8$ by seemingly minuscule modifications in the weights using $5\%$-$10\%$ of CIFAR10 samples (Figure 1). This result reflects that the robustness of UEs towards weight perturbations and uncertainties in classifier parameters is neither rigidly gauged nor carefully pondered.
@@
The aforementioned problems hinder the assessment of UEs and constitute additional risks to be exploited by deliberate adaptive adversaries. Therefore, it is crucial to develop a mechanism to ensure that a set of UEs can take effect without being ravaged by the problems of training stochasticity, cross-model generalization, and adaptive attackers. The mechanism can resort to finding classifiers that can achieve the best possible clean-task accuracy when trained on UEs and restraining this best-case accuracy. We are thus motivated to ask the question:
@@
[backgroundcolor=grey!10,rightline=true,leftline=true,topline=true,bottomline=true,roundcorner=1mm,everyline=true,nobreak=false] Is it possible to go beyond heuristics and provide an upper bound towards the clean-data performance of undivulged models trained on a set of UEs? The question is daunting to be thoroughly answered. However, we will show in this work that, it is possible to derive a clean-data accuracy upper bound guaranteed with a high probability, for unauthorized classifiers under some constraints. On the other hand, this threshold also serves as an upper bound of clean accuracy when facing the recovery attack. It can henceforth be a tool for gauging the effectiveness and robustness of UEs.
Our method and its broader impact.  In this paper, we propose a mechanism for certifying the best clean-task accuracy retained by unauthorized classifiers trained on a set of UEs. The mechanism can be applied to any empirically made UEs to derive a certified $(q,\eta)$-Learnability. $(q,\eta)$-Learnability gauges the effectiveness and robustness of UEs towards arbitrary classifiers sampled from a certain parameter space. Note that the certification pertains to the parameter space rather than the input space, alleviating the need for knowledge of the training strategies or learning algorithms used by the attackers. We also proposed certifying a generalization $(q,\eta)$-Learnability, which removes the requirement for holding private test datasets in the evaluation of UEs. To make the certification more sound, the paper also proposes a simple method to expand the certifiable parameter space and reduce the gap between the certified $(q,\eta)$-Learnability and the True Learnability (see Definition 1). Moreover, we produce Provably Unlearnable Examples (PUEs) which achieve lower $(q,\eta)$-Learnability scores and be more robust than existing PAP methods.
The $(q,\eta)$-Learnability metric serves as a rigid guarantee of data availability for machine learning models. This guarantee can operate in parallel with other techniques, such as deep watermarking [18] and differential privacy (DP) [19], to deter illegal data exploitation and support responsible AI regulations and standards (e.g., CCPA, GDPR, and Australia’s AI Safety Standard [20]). Instead of post-event ownership auditing through watermarks and protection of sensitive information with DP, $(q,\eta)$-Learnability proactively controls access to domain knowledge through data, safeguarding both data IP and privacy. The contributions of this paper are as follows:
@@
  • Our source code provides a tool for measuring the provable effectiveness of UEs and producing PUEs.
II Background
@@ II-A Unlearnable Examples
$\min_{\hat{f}}$ $\mathbb{E}_{(x,y)\sim{\mathbb{D}}_{s}}\min_{\delta}{\mathcal{L}}(\hat{f}(x+\delta),\,y)$ (2)
@@ II-B Smoothed Classifiers
$g(x)=$ $\operatorname*{arg\,max}_{y\in{\mathcal{Y}}}{\Pr_{\epsilon\sim\pi(x)}[f_{\theta}(x+\epsilon)=y]},\hskip 10.00002ptor$ (3)
@@
$\pi(x)$ is the distribution of the noise (i.e., smoothing distribution) centered at $x$ (or $\theta$) and $y$ is a predicted label from the label space ${\mathcal{Y}}$. The output from the smoothed classifier can be statistically bounded to offer certifiable robustness with respect to changes in the smoothed variables. Cohen et al. derived a tight $\ell_{2}$ robustness bound by smoothing classifier inputs with noise drawn from Gaussian distributions and certifying based on Neyman-Pearson Lemma [21]. Subsequently, the certified robustness is extended to various smoothing noises and certification frameworks [22, 23, 24, 25, 26, 27, 28, 29, 30]. Such smoothed classifier $g(x)$ can also be extended to randomized learning functions which in addition smooth the training dataset ${\mathbb{D}}_{s}$, resulting
$g({\mathbb{D}}_{s},x)=\operatorname*{arg\,max}_{y\in{\mathcal{Y}}}\Pr_{\hat{{\mathbb{D}}}_{s}\sim\pi({\mathbb{D}}_{s})}[f_{\theta}(\hat{{\mathbb{D}}}_{s},\ x)=y].$ (4)
@@
Figure 2: An overview of the certification and PUE crafting framework. A dataset ${\mathbb{D}}_{s}$ is perturbed into ${\mathbb{D}}_{s}\oplus\delta$ before being released to the public. The $(q,\eta)$-Learnability of ${\mathbb{D}}_{s}\oplus\delta$ can be certified to ensure that any unauthorized classifier trained on ${\mathbb{D}}_{s}\oplus\delta$ has a provable upper bound on its performance on any test set $\hat{{\mathbb{D}}}$ (or ${\mathbb{D}}$) in the same domain with ${\mathbb{D}}_{s}\oplus\delta$, as long as the parameters of the unauthorized classifier are within a certified parameter set. Generalization learnability can be computed using Hoeffding’s bound, and PAC-Bayesian theory suggests that a certification surrogate with low prediction variance under large parametric noise can improve certified learnability. Optimized PUEs lead to lower $(q,\eta)$-Learnability.
III Problem Statement
Problem definition.  Inheriting from the previous definitions, consider a to-be-published dataset ${\mathbb{D}}_{s}$ (e.g., portraits to be posted on Facebook) with $N$ data points. Additionally, there is a private dataset ${\mathbb{D}}$ (e.g., patient records with photos in a hospital database) sharing the same distribution ${\mathcal{D}}$ with ${\mathbb{D}}_{s}$. The defender’s goal is to render ${\mathbb{D}}_{s}$ unlearnable against unknown learning algorithms and classifiers (e.g., deep neural networks) by perturbing it into ${\mathbb{D}}_{s}\oplus\delta$, where $\delta$ is the anti-learning perturbation and $\oplus$ is the perturbing operator. We specifically consider additive perturbations $\delta=\{\delta_{i}\}^{|{\mathbb{D}}_{s}|}_{i=1}$ applied to each sample of ${\mathbb{D}}_{s}$ in this paper. Thus, $\oplus$ represents the element-wise addition. “Unlearnable” here means that a classifier trained on ${\mathbb{D}}_{s}\oplus\delta$ would exhibit a high classification error on ${\mathbb{D}}$. Furthermore, the defender wants to verify how robust ${\mathbb{D}}_{s}\oplus\delta$ is by having an upper bound on the accuracy that can be achieved by unknown classifiers stochastically trained on ${\mathbb{D}}_{s}\oplus\delta$. More profoundly, the defender wants to determine the best possible accuracy that classifiers trained on ${\mathbb{D}}_{s}\oplus\delta$ can achieve on ${\mathbb{D}}$. The defender does not desire an elevated best possible accuracy since it implies ${\mathbb{D}}_{s}\oplus\delta$ is exploitable. Therefore, the research goals are as follows. {mdframed}[backgroundcolor=grey!10,rightline=true,leftline=true,topline=true,bottomline=true,roundcorner=1mm,everyline=true,nobreak=false]
  • A certification mechanism should be developed to verify the maximum accuracy achievable by any classifiers and learning algorithms on ${\mathbb{D}}$ through training on ${\mathbb{D}}_{s}\oplus\delta$.
  • A better noise $\delta$ should be designed such that ${\mathbb{D}}_{s}\oplus\delta$ can effectively decrease the verified best accuracy.